# Accelerated Parallel Tempering via Neural Transports

### v3

@@ C.1.1 Work Formula for Deterministic flows
then, for all $(z^{n-1},z^{n})$ such that $T^{n}(z^{n-1})=z^{n}$, we have the following expression for the work defined in Equation˜3:
$W^{n}(z^{n-1},z^{n})=U^{n}(z^{n})-U^{n-1}(z^{n-1})-\log|\det J_{T^{n}}(z^{n-1})|;$
@@
$\frac{\mathrm{d}{\mathbb{Q}}^{n}}{\mathrm{d}{\mathbb{P}}^{n-1}}(z^{n-1},z^{n})=\frac{\pi^{n}(\mathrm{d}z^{n})}{(T^{n}\#\pi^{n-1})(\mathrm{d}z^{n})}\\ =\frac{\tilde{\pi}^{n}(z^{n})}{\tilde{\pi}^{n-1}((T^{n})^{-1}(z^{n}))|\det(J_{(T^{n})^{-1}}(z^{n}))|}=\frac{\tilde{\pi}^{n}(z^{n})|\det(J_{T^{n}}(z^{n-1}))|}{\tilde{\pi}^{n-1}(z^{n-1})}$ (32)
which justifies the identity for the work. Applying this at $(z^{n-1},z^{n})=(x^{n-1},x^{n}_{*})\in S$ gives
$\frac{\mathrm{d}{\mathbb{Q}}^{n}}{\mathrm{d}{\mathbb{P}}^{n-1}}(x^{n-1},x^{n}_{*})=\frac{\tilde{\pi}^{n}(x^{n}_{*})|\det(J_{T^{n}}(x^{n-1}))|}{\tilde{\pi}^{n-1}(x^{n-1})}.$ (33)
Similarly, applying Equation˜32 at $(z^{n-1},z^{n})=(x^{n-1}_{*},x^{n})\in S$ gives
$\frac{\mathrm{d}{\mathbb{Q}}^{n}}{\mathrm{d}{\mathbb{P}}^{n-1}}(x^{n-1}_{*},x^{n})=\frac{\tilde{\pi}^{n}(x^{n})|\det(J_{T^{n}}(x^{n-1}_{*}))|}{\tilde{\pi}^{n-1}(x^{n-1}_{*})}.$ (34)
Together Equation˜33 and Equation˜34 establish the proposition. ∎
C.1.2 Training
@@
There are also possible variants to the training pipeline. In each case, we initialize normalizing flows to the identity transformation. One option is to run the APT algorithm with the current flows. After every (or several) steps of the APT algorithm, the current samples can be used to update the flow parameters. Since the flows are initialized at the identity transformation, the initial sampling of APT behaves similarly to PT. A second possible training pipeline is to instead use PT directly to generate a large batch of samples, and then use this batch of samples to update the flows. The latter approach is more stable and is thus what we employed in the final experiments. We provide more details in Appendix˜D. We note that the first training pipeline has the potential to allow for better exploration, and we therefore leave a more detailed exploration of it to future work.
C.2 Further Details on Control Accelerated PT
@@
$W^{n}_{\infty}(x)$ $:=\Delta F_{n}-\log\frac{\mathrm{d}{\mathbb{Q}}^{n}_{\infty}}{\mathrm{d}{\mathbb{P}}^{n-1}_{\infty}}$
$=\int_{0}^{1}-\nabla\cdot b^{n}_{s}(x_{s})+\nabla U^{n}_{s}(x_{s})\cdot b^{n}_{s}(x_{s})+\partial_{s}U^{n}_{s}(x_{s})\mathrm{d}s,$
@@
$X_{0}\sim\pi^{n-1},\quad\mathrm{d}X_{s}=-(\sigma^{n}_{s})^{2}\nabla U^{n}_{s}(X_{s})\mathrm{d}s+b_{s}^{n}(X_{s})\mathrm{d}s+\sigma^{n}_{s}\sqrt{2}\mathrm{d}\accentset{\rightharpoonup}{B}_{s},$
@@
$X^{\prime}_{1}\sim\pi^{n},\quad\mathrm{d}X^{\prime}_{s}=(\sigma^{n}_{s})^{2}\nabla U^{n}_{s}(X^{\prime}_{s})\mathrm{d}s+b_{s}^{n}(X^{\prime}_{s})\mathrm{d}s+\sigma_{s}^{n}\sqrt{2}\mathrm{d}\accentset{\leftharpoonup}{B}_{s}.$
@@
C.3 Further Details on Diff-APT
@@
$dY_{s}=-\beta_{s}Y_{s}\mathrm{d}s+\sqrt{2\beta_{s}}\mathrm{d}W_{s},\quad Y_{0}\sim\pi,$
with the choice of schedule $\beta_{s}=\frac{1}{2(1-s)}$ to define the path of distributions $(\pi^{\text{VP}}_{s})_{s\in(0,1]}$ by $Y_{s}\sim\pi_{1-s}^{\text{VP}}$. Due to the singularity at $s=1$, $\pi_{0}^{\text{VP}}$ is not defined by the path, but we can define this point to be a standard Gaussian as the path converges to this distribution in the limit as $s$ approaches 1 in order to define the full annealing path $(\pi_{s}^{\text{VP}})_{s\in[0,1]}$.
@@
For a given annealing schedule $0=s_{0}<\ldots<s_{N}=1$, we construct $P_{k}^{n-1},Q_{k-1}^{n}$ through the linear discretisation of the time-reversal SDE on the time interval $s_{n-1},s_{n}]$ with step size $\delta_{n}=(s_{n}-s_{n-1})/K$ and interpolating times $s_{n,k}=s_{n-1}+k\delta_{n}$.
In particular, we take $Q_{k-1}^{n}(x_{k},dx_{k-1})$ to be ${\mathcal{N}}(\sqrt{1-\alpha_{n,k-1}}x_{k},\alpha_{n,k-1}\mathrm{I})$ where $\alpha_{n,k-1}=1-\exp(-2\int^{1-s_{n,k-1}}_{1-s_{n,k}}\gamma_{s}\mathrm{d}s)$ which is the closed-form kernel transporting $\pi_{s_{n,k}}^{\text{VP}}$ to $\pi_{s_{n,k-1}}^{\text{VP}}$. Furthermore, we take $P_{k}^{n-1}(x_{k-1},dx_{k})$ to be the exponential integrator given by ${\mathcal{N}}(\mu_{n,k-1}(x_{k-1}),\alpha_{n,k-1}\mathrm{I})$ where
$\mu_{n,k-1}(x)=\sqrt{1-\alpha_{n,k-1}}x+2(1-\sqrt{1-\alpha_{n,k-1}})(x+\nabla\log\pi_{s_{n,k-1}}^{\text{VP}}(x)).$
@@
$+$ $[1+r_{\theta}(1)-r_{\theta}(s)]\log g_{0}(\sqrt{s}x),$
where $g_{1}(x)\propto\pi(x)N(x;0,\mathrm{I})$. Here, $r$ is a scalar-valued neural network and $N$ is a vector-valued function in $\mathbb{R}^{d}$. We also note that $\pi^{\theta}_{1}(x)\propto\pi$, hence $\pi^{\theta}_{s}$ serves as a valid annealing path between $N(0,\mathrm{I})$ and $\pi$.
Appendix D Experimental Details
D.1 Target Distributions
@@
D.2 Network and Training Details
@@
For both GMM-$d$ and MW-32, we use a 4-layer MLP with 512 hidden units, and for DW-4, we use a 4-layer EGNN (Satorras et al., 2021) with 64 hidden units. Recall that in CMCD-APT, we use the linear path, $U^{\beta}(x)=(1-\beta)\log\eta(x)+\beta U(x)$, which linearly interpolates between reference and target in log-space. Therefore, our network is conditional on the $\beta$. We optimise the MLP by Adam with a learning rate of 1e-3 and EGNN with a learning rate of 1e-4. We additionally use gradient clipping with norm 1 for stability.
@@
  • Tuning PT: We define $\{\beta_{n}\}_{n=1}^{N}$ uniformly from 0 to 1. Then, we run PT for 600 steps, remove the first 100 steps as burn-in, and take the last 500 steps to calculate the rejection rate between adjacent chains. We then tune the value of $\beta_{n}$ to ensure the rejection rate even, according to Syed et al. (2021). We repeat this process 10 times to ensure $\{\beta_{n}\}_{n=1}^{N}$ is stable.
@@
For GMM-$d$, we use a 3-layer MLP with 128 hidden units, for MW-32, we use a 3-layer MLP with 256 hidden units and for DW-4, we use a 3 layer EGNN with 128 hidden units. For all models, we use a learning rate of 5e-4, gradient clipping with norm 1 for stability and EMA with decay parameter of 0.99. For training, we follow the same pipeline as CMCD-APT, but we only retain samples at the target chain in order to optimise the standard score matching objective. At sampling time, we tune the annealing schedule by running Diff-APT for 1,100 steps, discarding the first 100 samples as burn-in and using the last 1,000 steps to calculate rejection rates to apply the schedule tuning algorithm from Syed et al. (2021). We then repeat this 10 times where we initialise from the uniform schedule.
@@
D.3 Further Details on Comparison of Acceleration Methods
@@
D.4 Further Details on Scaling with Dimension
@@
D.5 Further Details on Free Energy Estimator
@@
D.6 Further Details on Comparing APT with Neural Samplers
For CMCD, we take our CMCD-APT model on DW-4 with 30 chains and $K=1,2,5$ from Section 6.3, and instead of sampling from these models using APT, we collect 5,000 independent samples from our reference distribution to which we apply our learned kernels $\prod_{n=1}^{N}\prod_{k=1}^{K}P_{k}^{n-1}$ which we recall are explicitly trained to transport samples from the reference to the target distribution. With our final samples, we then plot the histogram of $d_{ij}$ values for $i\neq j$.
For Diffusion, we follow the same procedure but we tune our annealing schedule with the same schedule tuning algorithm from Syed et al. (2021) as mentioned above (this step is not required for CMCD as the annealing schedule is required to be fixed during training) before we collect samples by applying $P_{k}^{n-1}$.
@@ D.6 Further Details on Comparing APT with Neural Samplers
D.7 Further Details on Alanine Dipeptide
We run CMCD-APT using 5 chains and $K=1,2,5$. For the local move, we adopt HMC with a dynamic step size: when the acceptance rate is larger than 0.9, we increase the step size by 1.2; if the acceptance rate is smaller than 0.8, we divide the step size by 1.2. Other settings are the same as those of other targets.
D.8 License
@@ D.9 Computing Resources
Appendix E Additional Experimental Results
E.1 Visualisation of ManyWell-32
@@
E.2 Comparison of Acceleration Methods for ManyWell-32 and DW-4
@@
Table 3: PT versus APT with different acceleration methods, targeting ManyWell-32 in 32 dimensions and standard Gaussian reference using $N=5,10,30$ parallel chains for $T=100,000$ iterations. For each method, we report the round trips (R), round trips per potential evaluation, denoted as compute-normalised round trips (CN-R), the number of neural network evaluations per parallel chain every iteration, and $\Lambda$ estimated using $N=30$ chains.
@@
Method Neural Call $(\downarrow)$ $\hat{\Lambda}$ $(\downarrow)$ R $(\uparrow)$ CN-R $(\uparrow)$ R $(\uparrow)$ CN-R $(\uparrow)$ R $(\uparrow)$ CN-R $(\uparrow)$
@@
Table 4: PT versus APT with different acceleration methods, targeting DW-4 in 10 dimensions and standard Gaussian reference using $N=5,10,30$ parallel chains for $T=100,000$ iterations. For each method, we report the round trips (R), round trips per potential evaluation, denoted as compute-normalised round trips (CN-R), the number of neural network evaluations per parallel chain every iteration, and $\Lambda$ estimated using $N=30$ chains.
@@
Method Neural Call $(\downarrow)$ $\hat{\Lambda}$ $(\downarrow)$ R $(\uparrow)$ CN-R $(\uparrow)$ R $(\uparrow)$ CN-R $(\uparrow)$ R $(\uparrow)$ CN-R $(\uparrow)$

### v4

@@ D.1.1 Work Formula for Deterministic flows
then, for all $(z^{n-1},z^{n})$ such that $T^{n}(z^{n-1})=z^{n}$, we have the following expression for the weight defined in Equation˜3:
$w^{n}(z^{n-1},z^{n})=\frac{\tilde{\pi}^{n}(z^{n})}{\tilde{\pi}^{n-1}(z^{n-1})}|\det J_{T^{n}}(z^{n-1})|$
@@
$\frac{\mathrm{d}{\mathbb{Q}}^{n}}{\mathrm{d}{\mathbb{P}}^{n-1}}(z^{n-1},z^{n})=\frac{\pi^{n}(\mathrm{d}z^{n})}{(T^{n}\#\pi^{n-1})(\mathrm{d}z^{n})}\\ =\frac{\pi^{n}(z^{n})}{\pi^{n-1}((T^{n})^{-1}(z^{n}))|\det(J_{(T^{n})^{-1}}(z^{n}))|}=\frac{\pi^{n}(z^{n})|\det(J_{T^{n}}(z^{n-1}))|}{\pi^{n-1}(z^{n-1})}$ (40)
which justifies the identity for the weight. Applying this at $(z^{n-1},z^{n})=(x^{n-1},x^{n}_{*})\in S$ gives
$\frac{\mathrm{d}{\mathbb{Q}}^{n}}{\mathrm{d}{\mathbb{P}}^{n-1}}(x^{n-1},x^{n}_{*})=\frac{\pi^{n}(x^{n}_{*})|\det(J_{T^{n}}(x^{n-1}))|}{\pi^{n-1}(x^{n-1})}.$ (41)
Similarly, applying Equation˜40 at $(z^{n-1},z^{n})=(x^{n-1}_{*},x^{n})\in S$ gives
$\frac{\mathrm{d}{\mathbb{Q}}^{n}}{\mathrm{d}{\mathbb{P}}^{n-1}}(x^{n-1}_{*},x^{n})=\frac{\pi^{n}(x^{n})|\det(J_{T^{n}}(x^{n-1}_{*}))|}{\pi^{n-1}(x^{n-1}_{*})}.$ (42)
Together Equation˜41 and Equation˜42 establish the proposition. ∎
D.1.2 Training
@@
There are also possible variants to the training pipeline. In each case, we initialize normalizing flows to the identity transformation. One option is to run the APT algorithm with the current flows. After every (or several) steps of the APT algorithm, the current samples can be used to update the flow parameters. Since the flows are initialized at the identity transformation, the initial sampling of APT behaves similarly to PT. A second possible training pipeline is to instead use PT directly to generate a large batch of samples, and then use this batch of samples to update the flows. The latter approach is more stable and is thus what we employed in the final experiments. We provide more details in Appendix E. We note that the first training pipeline has the potential to allow for better exploration, and we therefore leave a more detailed exploration of it to future work.
D.2 Further Details on Control Accelerated PT
@@
$W^{n}_{\infty}(x)$ $:=\log w^{n}_{\infty}(x)$
$=\int_{0}^{1}\nabla\cdot b^{n}_{s}(x_{s})+\nabla U^{n}_{s}(x_{s})\cdot b^{n}_{s}(x_{s})+\partial_{s}U^{n}_{s}(x_{s})\mathrm{d}s,$
@@
$X_{0}\sim\pi^{n-1},\quad\mathrm{d}X_{s}=(\sigma^{n}_{s})^{2}\nabla U^{n}_{s}(X_{s})\mathrm{d}s+b_{s}^{n}(X_{s})\mathrm{d}s+\sigma^{n}_{s}\sqrt{2}\mathrm{d}\accentset{\rightharpoonup}{B}_{s},$
@@
$X^{\prime}_{1}\sim\pi^{n},\quad\mathrm{d}X^{\prime}_{s}=-(\sigma^{n}_{s})^{2}\nabla U^{n}_{s}(X^{\prime}_{s})\mathrm{d}s+b_{s}^{n}(X^{\prime}_{s})\mathrm{d}s+\sigma_{s}^{n}\sqrt{2}\mathrm{d}\accentset{\leftharpoonup}{B}_{s}.$
@@
D.3 Further Details on Diff-APT
@@
$dY_{s}=-\gamma_{s}Y_{s}\mathrm{d}s+\sqrt{2\gamma_{s}}\mathrm{d}W_{s},\quad Y_{0}\sim\pi,$
with the choice of schedule $\gamma_{s}=\frac{1}{2(1-s)}$ to define the path of distributions $(\pi^{\text{VP}}_{s})_{s\in(0,1]}$ by $Y_{s}\sim\pi_{1-s}^{\text{VP}}$. Due to the singularity at $s=1$, $\pi_{0}^{\text{VP}}$ is not defined by the path, but we can define this point to be a standard Gaussian as the path converges to this distribution in the limit as $s$ approaches 1 in order to define the full annealing path $(\pi_{s}^{\text{VP}})_{s\in[0,1]}$.
@@
For a given annealing schedule $0=s_{0}<\ldots<s_{N}=1$, we construct $P_{k}^{n-1},Q_{k-1}^{n}$ through the linear discretisation of the time-reversal SDE on the time interval $s_{n-1},s_{n}]$ with step size $\delta_{n}=(s_{n}-s_{n-1})/K$ and interpolating times $s_{n}^{k}=s_{n-1}+k\delta_{n}$.
In particular, we take $Q_{k-1}^{n}(x_{k},dx_{k-1})$ to be ${\mathcal{N}}(\sqrt{1-\alpha_{n,k-1}}x_{k},\alpha_{n,k-1}\mathrm{I})$ where $\alpha_{n,k-1}=1-\exp(-2\int^{1-s_{n}^{k-1}}_{1-s_{n}^{k}}\gamma_{s}\mathrm{d}s)$ which is the closed-form kernel transporting $\pi_{s_{n}^{k}}^{\text{VP}}$ to $\pi_{s_{n}^{k-1}}^{\text{VP}}$. Furthermore, we take $P_{k}^{n-1}(x_{k-1},dx_{k})$ to be the exponential integrator given by ${\mathcal{N}}(\mu_{n,k-1}(x_{k-1}),\alpha_{n,k-1}\mathrm{I})$ where
$\mu_{n,k-1}(x)=\sqrt{1-\alpha_{n,k-1}}x+2(1-\sqrt{1-\alpha_{n,k-1}})(x+\nabla\log\pi_{s_{n}^{k-1}}^{\text{VP}}(x)).$
@@
$+$ $[1+r_{\theta}(1)-r_{\theta}(s)]\log g_{1}(\sqrt{s}x),$
where $\pi(x)\propto g_{1}(x)N(x;0,\mathbf{I})$. Here, $r$ is a scalar-valued neural network and $N$ is a vector-valued function in $\mathbb{R}^{d}$. We also note that $\pi^{\theta}_{1}(x)\propto\pi$, hence $\pi^{\theta}_{s}$ serves as a valid annealing path between $N(0,\mathbf{I})$ and $\pi$.
Appendix E Experimental Details
E.1 Target Distributions
@@
E.2 Network and Training Details
@@
For both GMM-$d$ and MW-32, we use a 4-layer MLP with 512 hidden units, and for DW-4, we use a 4-layer EGNN (Satorras et al., 2021) with 64 hidden units. Recall that in CMCD-APT, we use the geometric path, $\pi^{\beta}(x)=\eta(x)^{1-\beta}\pi(x)^{\beta}$, which linearly interpolates between reference and target in log-space. Therefore, our network is conditional on the $\beta$. We optimise the MLP by Adam with a learning rate of 1e-3 and EGNN with a learning rate of 1e-4. We additionally use gradient clipping with norm 1 for stability.
@@
  • Tuning PT: We initialise $\{\beta_{n}\}_{n=1}^{N}$ uniformly from 0 to 1. Then, we run PT for 600 steps, remove the first 100 steps as burn-in, and take the last 500 steps to calculate the rejection rate between adjacent chains using this to apply Algorithm˜2 to update the schedule. We repeat this process 10 times to ensure $\{\beta_{n}\}_{n=1}^{N}$ is stable.
@@
For GMM-$d$, we use a 3-layer MLP with 128 hidden units, for MW-32, we use a 3-layer MLP with 256 hidden units and for DW-4, we use a 3 layer EGNN with 128 hidden units. For all models, we use a learning rate of 5e-4, gradient clipping with norm 1 for stability and EMA with decay parameter of 0.99. For training, we follow the same pipeline as CMCD-APT, but we only retain samples at the target chain in order to optimise the standard score matching objective. At sampling time, we tune the annealing schedule by running Diff-APT for 1,100 steps, discarding the first 100 samples as burn-in and using the last 1,000 steps to calculate rejection rates. We then use this to apply Algorithm˜2 to update the schedule. We then repeat this 10 times where we initialise from the uniform schedule.
@@
E.3 Further Details on Comparison of Acceleration Methods
@@
E.4 Further Details on Scaling with Dimension
@@
E.5 Further Details on Log-Normalising Constant (Free-Energy) Estimation
@@
E.6 Further Details on Comparing APT with Neural Samplers
For CMCD, we take our CMCD-APT model on DW-4 with 30 chains and $K=1,2,5$ from Section 6.3, and instead of sampling from these models using APT, we collect 5,000 independent samples from our reference distribution to which we apply our learned kernels $\prod_{n=1}^{N}\prod_{k=1}^{K}P_{k}^{n-1}$ which we recall are explicitly trained to transport samples from the reference to the target distribution. With our final samples, we then plot the histogram of $d_{ij}$ values for $i\neq j$—specifically, we note that DW-4 represents a distribution over 4 particles in 2D, therefore for each sample from DW-4, we compute the pairwise distances $d_{ij}$ ($i\neq j$) over these particles in the sample; we then collect all of these values to plot their histogram.
For Diffusion, we follow the same procedure but we tune our annealing schedule with Algorithm˜2 applied in the same way as in Appendix E.2 (this step is not required for CMCD as the annealing schedule is required to be fixed during training) before we collect samples by applying $P_{k}^{n-1}$.
@@ E.6 Further Details on Comparing APT with Neural Samplers
E.7 Further Details on Alanine Dipeptide
We run CMCD-APT using 4 chains and $K=1,2,5$. For the local move, we adopt HMC with a dynamic step size: when the acceptance rate is larger than 0.9, we increase the step size by 1.2; if the acceptance rate is smaller than 0.8, we divide the step size by 1.2. Other settings are the same as those of other targets.
E.8 License
@@ E.9 Computing Resources
Appendix F Additional Experimental Results
F.1 Visualisation of ManyWell-32
@@
F.2 Comparison of Acceleration Methods for ManyWell-32 and DW-4
@@
Table 3: PT versus APT with different acceleration methods, targeting ManyWell-32 in 32 dimensions and standard Gaussian reference using $N=5,10,30$ parallel chains for $T=100,000$ iterations. For each method, we report the round trips (R), round trips per target evaluation, denoted as compute-normalised round trips (CN-R), the number of neural network evaluations per parallel chain every iteration (Neural Calls), and $\Lambda_{K}$ estimated using $N=30$ chains $(\hat{\Lambda}_{K})$.
@@
Method Neural Calls $(\downarrow)$ $\hat{\Lambda}_{K}$ $(\downarrow)$ R $(\uparrow)$ CN-R $(\uparrow)$ R $(\uparrow)$ CN-R $(\uparrow)$ R $(\uparrow)$ CN-R $(\uparrow)$
@@
Table 4: PT versus APT with different acceleration methods, targeting DW-4 in 8 dimensions and standard Gaussian reference using $N=5,10,30$ parallel chains for $T=100,000$ iterations. For each method, we report the round trips (R), round trips per target evaluation, denoted as compute-normalised round trips (CN-R), the number of neural network evaluations per parallel chain every iteration (Neural Calls), and $\Lambda_{K}$ estimated using $N=30$ chains $(\hat{\Lambda}_{K})$.
@@
Method Neural Calls $(\downarrow)$ $\hat{\Lambda}_{K}$ $(\downarrow)$ R $(\uparrow)$ CN-R $(\uparrow)$ R $(\uparrow)$ CN-R $(\uparrow)$ R $(\uparrow)$ CN-R $(\uparrow)$